The numerator and denominator of a fraction are in ratio 2 : 3. If 6 is subtracted from the numerator, the
result will be 2/3 of the original fraction. What is the value of the numerator ?

Answer by mananth(16946)

x/y = 2/3
3x=2y
3x-2y=0............(1)

(x-6)/y = (2/3) (x/y)
3x-18 = 2x
x=18